40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Apr 7 23:18:49 2011, koji, frank, tara, DailyProgress, Seismic, previous suspension for RCAV restored 000_0014.jpg000_0018.jpg
    Reply  Sat Apr 9 07:11:20 2011, koji, frank, tara, DailyProgress, Seismic, Q measurement 6x
Message ID: 571     Entry time: Sat Apr 9 07:11:20 2011     In reply to: 569
Author: koji, frank, tara 
Type: DailyProgress 
Category: Seismic 
Subject: Q measurement 

Q measurement from the cavity's suspension is measured. Vertical translational mode is 3.46 Hz, Q=52 (previous setup was ~200 .) Pitch mode is damped very fast and turns into vertical mode quickly. Q measurement for new periscope is measured and compared with one of the current periscopes (the one for exiting beam, RCAV). The new one is quite better.

 

The suspension of RCAV is restored, with better eddy current damping. We used shadow sensing technique to measured Q of vertical translational motion of the cavity.

f0 = 3.46 Hz, Q = 45.

cavity_v.png

However, pitch mode dies down very fast and turns into vertical mode. We could not really measure it properly. The data looks like this.

pitch.png

 

Next, we measured Q factors for the new periscope that will replace the current one. The current one is made by adding a few 2"-3" posts together, the mirror mounts

are also shaky. So before replacing it, we want to see how better the new one will be.

For the new periscope, there will be one mirror mounted on the post. The bottom mirror will be mounted on the table. So the setup is shown below.

HeNe beam reflects at the mirror on the mirror mount. The post is tapped on two directions, pitch and yaw. They are slightly different due to asymmetry introduced by the top mirror mount.

Then the beam falls on the PD, which is mounted on the table to reduce any extra resonant peaks.

The beam is clipped by a razor blade attached on the PD when the beam moves.

We also increased the sensitivity by focusing the beam to a smaller spot.

Qmsmt.jpg

First, we measured the PSD of the PD signal by tapping on each directions because we could not excite it with white noise, and we wanted to see the peak from each mode clearly.

Once we see the resonance peak, we can fit it to see the Q of each peak

 newperiscope.png

Mode    Resonant freq    Q

yaw          221                27

pitch        249.5               70

 Note: the peak at 108 Hz is the resonant frequency of the laser mount.

The result from ring down measurement is quite different, the pitch frequency is the same, but Q is 45.

The discrepancy might come from the fact that when we measured the ring down, we did not excite exactly one mode, so we see the result from the combination of pitch and yaw.  But the number is not outrageously different anyway.

newperi_ringdown.png

 

Then we tested the curernt periscope. The setup was quite similar except we removed the bottom mirror, so any peaks from bottom mirror will not be seen here

and its peaks' resonant frequencies might shift a bit due to the missing mirror' mass. The results are

 

Mode    Res freq[Hz]      Q

               131          120

              148           100

              526            40

              886            30

 

badperi.png

I could not identify which modes cause which frequencies. But those peaks do not appear in the new periscope test, so they are likely to be from the bad one.

I'm surprised that those measured peaks do not appear in the beat measurement.

 

 

ELOG V3.1.3-